# Beyond Simple Graphs: Neural Multi-Objective Routing on Multigraphs

### v4

@@ Abstract
Learning-based methods for routing have gained significant attention in recent years, both in single-objective and multi-objective contexts. Yet, existing methods are unsuitable for routing on multigraphs, which feature multiple edges with distinct attributes between node pairs, despite their strong relevance in real-world scenarios. In this paper, we propose two graph neural network-based methods to address multi-objective routing on multigraphs. Our first approach operates directly on the multigraph by autoregressively selecting edges until a tour is completed. The second model, which is more scalable, first simplifies the multigraph via a learned pruning strategy and then performs autoregressive routing on the resulting simple graph. We evaluate both models empirically, across a wide range of problems and graph distributions, and demonstrate their competitive performance compared to strong heuristics and neural baselines [^1] Our code will be released publicly upon paper acceptance..
@@ 1 Introduction
In spite of the high practical relevance of multigraph formulations (Lai et al., 2016; Ben Ticha et al., 2017), current learning-based methods are incapable of handling them due to two main reasons. Firstly, many state-of-the-art neural solvers rely on transformers to encode the problem instance. While these work well in the Euclidean setting (Kool et al., 2018) and with some modifications on asymmetric, directed graphs (Kwon et al., 2021), they lack the capability to encode multigraph structures. Secondly and more importantly, planning routes in multigraphs requires both selecting the node order and which edges to traverse, making current decoding strategies unsuitable.
@@
  • We design models to handle multigraph inputs. To our knowledge, we present the first neural solvers designed for such structures. In Appendix A, we argue further why current methods are insufficient, even when combined with graph transformations.
  • Our models are also capable of handling a variety of MO routing problems on asymmetric graphs. In contrast, current neural MO methods are either problem-specific or designed exclusively for the Euclidean setting, where edge costs are symmetric and determined only by the node coordinates. The non-Euclidean, asymmetric setting is generally viewed as more practically relevant (Boyacı et al., 2021).
@@ 2.1 Learning for Routing
Motivated by the success of end-to-end learning in fields such as image classification and language modeling, neural solutions for routing have gained increasing attention during the last years. Broadly, methods can be classified by how they obtain tours. Two prominent alternatives are autoregressive construction, based on iterative building of tours one node at a time, and non-autoregressive construction, that outputs heatmaps in a one-shot fashion for down-stream decoding with some search method. Examples from the first category include Vinyals et al. (2015); Kool et al. (2018); Kwon et al. (2020). Examples from the second category are Joshi et al. (2019); Fu et al. (2021); Sun & Yang (2023). For a more comprehensive discussion of these methods, we refer the reader to Appendix B, which presents additional related work.
@@ 3 Problem Formulation
Given a solution in the PS, there always exists a preference $\lambda$ and corresponding Chebyshev subproblem for which the solution is optimal (Choo & Atkins, 1983).
@@ 4.1 Edge-Based GMS
Here, $\theta(\lambda)$ is the weights of the model given the current preference $\lambda$. As Navon et al. (2021) and Lin et al. (2022), we utilize a Multi-Layer Perceptron (MLP) hyper-network to preference-condition the decoder, while keeping the encoder preference-agnostic. Consequently we form the decoder weights according to $\theta_{\text{dec}}(\lambda)=\text{MLP}(\lambda)$ and set $\theta(\lambda)=[\theta_{\text{enc}},\theta_{\text{dec}}(\lambda)]$. Regarding the decoder architecture, it is an edge-based variant of the Multi-Pointer (MP) decoder from Jin et al. (2023). We detail this component further in Appendix C.
[…]
Figure 1: Edge-based GMS and its most important components.
Note that a multigraph with $N$ nodes and $M$ edges between each pair of nodes has $\mathcal{O}(MN^{2})$ edges in total. Moreover, the decoder must be run $\mathcal{O}(N)$ times to construct a full tour and we roll out $\mathcal{O}(N)$ sample trajectories according to the POMO framework (Kwon et al., 2020). Hence, an edge-based decoder scales as $\mathcal{O}(MN^{4})$ compared to a node-based one that scales as $\mathcal{O}(N^{3})$. Consequently, we propose a second architecture that employs a node-based decoder and thereby scales better to higher node counts in terms of memory requirements and runtime.
@@ 4.2 Dual Head GMS
From a broader perspective, our approach in GMS-DH can be viewed as a combination of non-autoregressive and autoregressive construction. As a drawback, the edge choice can no longer depend on the context given by the partial tour. Instead, it must be chosen before starting roll-out.
@@ 4.3 Training Algorithm
For the second model, the training is more involved. Given an edge selection $\mathcal{E}\sim q_{\tilde{\theta}(\lambda)}$ from the selection-head according to equation 10, the routing-head parameterizes a policy $p_{\theta(\lambda)}(\cdot\mid s,\mathcal{E})$ and obtains a scalar reward $R^{(2)}_{\lambda}(\pi)=R_{\lambda}(\pi)$ by equation 11 as usual. With a fixed $\tilde{\theta}$, the objective for this head is to maximize
@@ 5 Experimental Studies
In the multigraph setting, we consider the Multigraph MOTSP (MGMOTSP) with two objectives, where we sample two-dimensional edge distances using distributions called FIXxx and FLEXxx. Here, $x$ denotes the number of edges between each node pair. In the FLEX-distribution, we sample edge distances independently and remove edges which are dominated. In the FIX distribution we first sample edges independently, but then rearrange by sorting, so that the an edge that is best in one objective is the worst in the other objective. Thus, in the FIXxx-distribution there are always exactly $x$ edges between each node pair, whereas FLEXxx has at most $x$ edges.
@@
Baselines. For the MOTSP and MGMOTSP, we utilize linear scalarization together with strong single-objective methods in the form of LKH3 (Helsgaun, 2017) and Google OR-tools (Perron & Furnon, 2019). We also benchmark against four Multi-Objective Evolutionary Algorithms (MOEAs): MOGLS (Ishibuchi & Murata, 1998), MOEA/D (Zhang & Li, 2007), NSGA-II (Deb et al., 2002) and NSGA-III (Blank et al., 2019). Below, we only show the results for the MOEA that performs best in each scenario.
Finally, we also design a neural baseline based on the MatNet-architecture (Kwon et al., 2021) as encoder and a preference-conditioned MP network as decoder. We call this MatNet-Based Model (MBM) and our motivation is that it is a method that can be used almost ”off-the-shelf” by combining components from literature. In the multigraph setting, this architecture requires us to pre-process the input by pruning edges until we obtain a simple graph, which we do using linear scalarization following Proposition 1.
@@
Model Settings. We train all models for $200$ epochs using 100 000100\,000 randomly generated instances per epoch with the ADAM optimizer (Kingma & Ba, 2015). We use the batch size $B=64$ and the learning-rate $\eta=10^{-4}$. For each batch, we sample a preference $\lambda=(\lambda_{1},\lambda_{2})$ using $\lambda_{1}\sim\text{Unif}[0,1]$ and $\lambda_{2}=1-\lambda_{1}$. For GMS-EB we let the number of layers be $L=6$, while GMS-DH has $L_{1}=5$ and $L_{2}=2$. Both models have embedding dimension 128 and use $8$ attention heads in the GREAT layers and MP decoders. Regarding the number of samples in the decoding, we set $K_{1}=4$ for the selection-decoder during training and $K_{1}=1$ during inference. In the routing-decoder, we always set $K_{2}$ to the problem size as customary in the POMO framework (Kwon et al., 2020). Finally, we augment the models during inference with a factor of $8$. For MBM, this is done using the original MatNet-augmentation (Kwon et al., 2021), while for GMS we use the scaling-augmentation of Lischka et al. (2025).
@@ 5.1 Performance on Benchmark Problems
We show the performance of GMS together with the other methods in Table 1. Both variants of GMS outperform most benchmarks on all problems. On the MOCVRP they are the best performing methods, while only LKH outperform them on the MOTSP and MGMOTSP. This showcases that our architectures are generally applicable across a variety of distributions, instance sizes and problems. As for the baseline neural method, MBM, it performs quite well across many cases, especially on the MGMOTSP, but severely under performs on some distributions, e.g., XASY100 for the MOCVRP.
Table 1: MOTSP, MOCVRP and MGMOTSP with different instance sizes (50, 100) and distributions. Runtime is the total time for solving 200 instances. The best results (excluding LKH, our baseline) are in bold and the second best are underlined. Remark that most neural MO solvers benchmark on instances of roughly this size (Chen et al., 2023b; Fan et al., 2025).
@@
However, MBM is usually slightly faster than GMS-DH, owing to the MatNet architecture being faster and more memory-efficient than GREAT (for a comparison, see the results by Lischka et al. (2025)). GMS-DH has a small advantage for the MGMOTSP with $50$ nodes due to only encoding each multigraph once, but this is offset again for $100$ nodes. Nonetheless, both GMS models remain efficient relative to the other benchmarks, despite the slower scaling of GMS-EB. In particular, GMS-DH as well as GMS-EB without augmentation are significantly faster than LKH.
Apart from the previously mentioned benchmarks, we also evaluate a variant of GMS-DH with Pre-Pruning (GMS-DH PP) applied before the encoder in the MGMOTSP setting. This pre-pruning is the same as for MBM, and the selection head is removed accordingly.
@@ 5.2 Zero-Shot Generalization
We also test our models on larger instances not seen during training, and report the results in Table 3. Both GMS-DH and GMS-EB remain competitive across both problems and all distributions in this zero-shot setting.
@@
5.3 A Difficult Multigraph Problem
Finally, we revisit one of the main motivations behind this work: routing for problems without clear a priori edge selection. Thus, we evaluate the performance on a more complex problem. Specifically, inspired by Ben Ticha et al. (2017), we consider an MGMOTSP with Time-Windows (MGMOTSPTW). Each edge is associated with a time and distance, while the objectives correspond to the number of violated time-windows and the total distance. Remark that, unlike the MGMOTSP, it is not clear in this case which edges can be pruned due to the more intricate time-window objective.
@@ 6 Conclusion
In this paper, we propose two learning-based approaches for solving multi-objective routing problems, such as the MOTSP, on multigraphs and asymmetric graphs. Both methods utilize graph neural networks for encoding the graph structure but differ in their route construction strategies. The first method is edge-based and directly builds tours on the multigraph, while the second employs a two-stage decoding process, consisting of non-autoregressive pruning of the multigraph followed by route construction.
Empirical results show that both methods perform well across a variety of settings. While the edge-based model is simpler and usually demonstrates slightly superior performance, the second model is more scalable due to lower memory and runtime demands.
@@ Acknowledgements
Appendix A Extended Motivation
In this section, we further motivate our approach by explaining why simple transformations of the underlying multigraph, combined with vanilla node-based construction, are unsuitable in the context of vehicle routing problems.

### v5

@@ Abstract
Learning-based methods for routing have gained significant attention in recent years, both in single-objective and multi-objective contexts. Yet, existing methods are unsuitable for routing on multigraphs, which feature multiple edges with distinct attributes between node pairs, despite their strong relevance in real-world scenarios. In this paper, we propose two graph neural network-based methods to address multi-objective routing on multigraphs. Our first approach operates directly on the multigraph by autoregressively selecting edges until a tour is completed. The second model, which is more scalable, first simplifies the multigraph via a learned pruning strategy and then performs autoregressive routing on the resulting simple graph. We evaluate both models empirically, across a wide range of problems and graph distributions, and demonstrate their competitive performance compared to strong heuristics and neural baselines******Our implementation is available at https://github.com/filiprydin/GMS..
@@ 1 Introduction
In spite of the high practical relevance of multigraph formulations (Lai et al., 2016; Ben Ticha et al., 2017), current learning-based methods are not equipped to handle them due to two main reasons. Firstly, many state-of-the-art neural solvers rely on transformers to encode the problem instance. While these work well in the Euclidean setting (Kool et al., 2018) and with some modifications on asymmetric, directed graphs (Kwon et al., 2021), they are ill-suited for encoding multigraph structures. Secondly and more importantly, planning routes in multigraphs requires both selecting the node order and which edges to traverse, making most current decoding strategies unsuitable.
@@
  • We design models to handle multigraph inputs. To our knowledge, we present the first neural solvers specifically designed for such structures. In Appendix B, we argue further why current methods are insufficient, even when combined with graph transformations.
  • Our models are also capable of handling a variety of MO routing problems on asymmetric graphs. In contrast, current neural multi-objective methods are either problem-specific or designed exclusively for the Euclidean setting, where edge costs are symmetric and determined only by the node coordinates. The non-Euclidean, asymmetric setting is generally viewed as more practically relevant (Boyacı et al., 2021). In our results, we show that naive extensions of many single-objective neural methods for asymmetric problems fail to perform well, illustrating the difficulty of the multi-objective non-Euclidean setting.
@@ 2.1 Learning for Routing
Motivated by the success of end-to-end learning in fields such as image classification and language modeling, neural solutions for routing have gained increasing attention during the last years. Broadly, methods can be classified by how they obtain tours. Two prominent alternatives are autoregressive construction, based on iterative building of tours one node at a time, and non-autoregressive construction, that outputs heatmaps in a one-shot fashion for down-stream decoding with some search method. Examples from the first category include Vinyals et al. (2015); Kool et al. (2018); Kwon et al. (2020). Examples from the second category are Joshi et al. (2019); Fu et al. (2021); Sun and Yang (2023). For a more comprehensive discussion of these methods, we refer the reader to Appendix A, which presents additional related work.
@@ 3 Problem Formulation
Given a solution in the PS, there always exists a preference $\lambda$ and corresponding Chebyshev subproblem for which the solution is optimal (Choo and Atkins, 1983).
@@ 4.1 Edge-Based GMS
Here, $\theta(\lambda)$ is the weights of the model given the current preference $\lambda$. As Navon et al. (2021) and Lin et al. (2022), we utilize a Multi-Layer Perceptron (MLP) hyper-network to preference-condition the decoder, while keeping the encoder preference-agnostic. Consequently we form the decoder weights according to $\theta_{\text{dec}}(\lambda)=\text{MLP}(\lambda)$ and set $\theta(\lambda)=[\theta_{\text{enc}},\theta_{\text{dec}}(\lambda)]$. Regarding the decoder architecture, it is an edge-based variant of the Multi-Pointer (MP) decoder from Jin et al. (2023). We detail this novel component further in Appendix C.
Note that a multigraph with $N$ nodes and $M$ edges between each pair of nodes has $\mathcal{O}(MN^{2})$ edges in total. In our decoding, we only calculate scores for outgoing edges from each node, but all nodes will eventually be visited and thus all edge embeddings must be stored in memory. Moreover, the decoder must be run $\mathcal{O}(N)$ times to construct a full tour and we roll out $\mathcal{O}(N)$ sample trajectories according to the POMO framework (Kwon et al., 2020). Hence, an edge-based decoder scales as $\mathcal{O}(MN^{4})$ compared to a node-based one that scales as $\mathcal{O}(N^{3})$. Consequently, we propose a second architecture that employs a node-based decoder and thereby scales better to higher node counts in terms of memory requirements and runtime.
@@ 4.2 Dual Head GMS
Figure 1: Edge-based GMS and its most important components.
@@
From a broader perspective, our approach in GMS-DH can be viewed as a combination of Autoregressive (AR) and Non-Autoregressive (NAR) construction. As a drawback, the edge choice can no longer depend on the context given by the partial tour. Instead, it must be chosen before starting roll-out. Combinations of AR and NAR construction have previously been explored for divide-and-conquer approaches to tackle large-scale VRPs (Ye et al., 2024). However, this combination of NAR pruning and AR construction, and using a single joint encoder, is novel and specifically designed for the multigraph setting. We further clarify the technical novelty of the different components in GMS-DH, as well as GMS-EB, in Appendix B.3.
@@ 4.3 Training Algorithm
For the second model, we propose a more involved training setup. Given an edge selection $\mathcal{E}\sim q_{\tilde{\theta}(\lambda)}$ from the selection-head according to equation 10, the routing-head parameterizes a policy $p_{\theta(\lambda)}(\cdot\mid s,\mathcal{E})$ and obtains a scalar reward $R^{(2)}_{\lambda}(\pi)=R_{\lambda}(\pi)$ by equation 11 as usual. With a fixed $\tilde{\theta}$, the objective for this head is to maximize
@@ 5 Experimental Studies
In the multigraph setting, we consider the Multigraph MOTSP (MGMOTSP) and Multigraph MOCVRP (MGMOCVRP) with two objectives, where we sample two-dimensional edge distances using distributions called FIXxx and FLEXxx. Here, $x$ denotes the number of edges between each node pair. In the FLEX-distribution, we sample edge distances independently and remove edges which are dominated. In the FIX distribution we first sample edges independently, but then rearrange by sorting, so that the an edge that is best in one objective is the worst in the other objective. Thus, in the FIXxx-distribution there are always exactly $x$ edges between each node pair, whereas FLEXxx has at most $x$ edges.
@@
Baselines. For the MOTSP, MGMOTSP and MGMOCVRP, we utilize linear scalarization together with strong single-objective methods in the form of LKH3 (Helsgaun, 2017), Google OR-tools (Perron and Furnon, 2019) and HGS (Vidal et al., 2012). We also benchmark against four Multi-Objective Evolutionary Algorithms (MOEAs): MOGLS (Ishibuchi and Murata, 1998), MOEA/D (Zhang and Li, 2007), NSGA-II (Deb et al., 2002) and NSGA-III (Blank et al., 2019). Below, we only show the results for the MOEA that performs best in each scenario.
Finally, we also design a neural baseline based on the MatNet-architecture (Kwon et al., 2021) as encoder and a preference-conditioned MP network as decoder. We call this MatNet-Based Model (MBM) and our motivation is that it is a method that can be used almost “off-the-shelf” by combining components from literature. In the multigraph setting, this architecture requires us to pre-process the input by pruning edges until we obtain a simple graph, which we do using linear scalarization following Proposition 1.
@@
Model Settings. We train all models for $200$ epochs using 100 000100\,000 randomly generated instances per epoch with the ADAM optimizer (Kingma and Ba, 2015). We use the batch size $B=64$ and the learning-rate $\eta=10^{-4}$. For each batch, we sample a preference $\lambda=(\lambda_{1},\lambda_{2})$ using $\lambda_{1}\sim\text{Unif}[0,1]$ and $\lambda_{2}=1-\lambda_{1}$. For GMS-EB we let the number of layers be $L=6$, while GMS-DH has $L_{1}=5$ and $L_{2}=2$. Both models have embedding dimension 128 and use $8$ attention heads in the GREAT layers and MP decoders. Regarding the number of samples in the decoding, we set $K_{1}=4$ for the selection-decoder during training and $K_{1}=1$ during inference. In the routing-decoder, we always set $K_{2}$ to the problem size as customary in the POMO framework (Kwon et al., 2020). Finally, we augment the models during inference with a factor of $8$. For MBM, this is done using the original MatNet-augmentation (Kwon et al., 2021), while for GMS we use the scaling-augmentation of Lischka et al. (2025).
@@ 5.1 Performance on Benchmark Problems
Table 1: MOTSP, MOCVRP, MGMOTSP and MGMOCVRP with different instance sizes (50, 100) and distributions. Runtime is the total time for solving 200 instances. The best results (excluding LKH and HGS, our baselines) are in bold and the second best are underlined. Remark that most neural MO solvers benchmark on instances of roughly this size (Chen et al., 2023b; Fan et al., 2025a).
@@
FLEX2-50 FLEX2-100 FIX2-50 FIX2-100
MGMOCVRP HGS 0.89 0.00% (15h) 0.89 0.00% (75h) 0.87 0.00% (15h) 0.91 0.00% (74h)
OR Tools 0.84 5.93% (1.5h) 0.84 5.97% (1.9h) 0.83 5.14% (1.5h) 0.88 3.78% (1.9h)
NSGA-II 0.71 20.05% (37h) 0.56 36.54% (219h) 0.65 25.16% (36h) 0.61 33.28% (224h)
MBM 0.83 6.07% (16s) 0.81 9.38% (54s) 0.84 4.14% (16s) 0.87 5.00% (54s)
MBM (aug) 0.84 5.23% (2.0m) 0.82 8.30% (6.9m) 0.84 3.39% (1.9m) 0.87 4.32% (6.9m)
GMS-EB 0.85 4.12% (1.2m) 0.85 4.08% (10m) 0.84 3.47% (1.2m) 0.89 2.81% (9.9m)
GMS-EB (aug) 0.86 3.55% (9.0m) 0.86 3.66% (1.4h) 0.85 2.71% (11m) 0.89 2.38% (1.4h)
GMS-DH 0.85 4.86% (20s) 0.85 4.94% (1.0m) 0.83 4.96% (20s) 0.88 3.58% (1.0m)
GMS-DH (aug) 0.85 3.86% (2.1m) 0.85 3.89% (8.2m) 0.84 3.96% (2.1m) 0.88 3.05% (8.1m)
GMS-DH PP 0.75 15.60% (36s) 0.75 15.70% (2.45m) 0.76 12.82% (36s) 0.81 11.05% (2.45m)
We show the performance of GMS together with the other methods in Table 1. Both variants of GMS outperform most benchmarks on all problems. On the MOCVRP they are the best performing methods, while only LKH outperform them on the MOTSP and MGMOTSP. For the MGMOCVRP, HGS attains the highest hypervolume, with GMS-EB and GMS-DH following in all cases except FIX2-50, where MBM performs better than GMS-DH. This showcases that our architectures are effective across a variety of distributions, instance sizes and problems. The baseline neural method, MBM, also performs quite well across many cases, especially on the MGMOTSP and MGMOCVRP, but severely underperforms on some distributions, e.g., XASY100 for the MOCVRP. In Appendix H, we include results for other naive extensions of existing neural methods, and show that they also tend to be less effective and robust compared to GMS-EB and GMS-DH.
However, MBM is usually slightly faster than GMS-DH, owing to the MatNet architecture being faster and more memory-efficient than GREAT (for a direct comparison, see the results by Lischka et al. (2025)). GMS-DH has a small advantage for the MGMOTSP with $50$ nodes due to only encoding each multigraph once, but this is offset again for $100$ nodes. Nonetheless, both GMS models remain efficient relative to the other benchmarks, despite the slower scaling of GMS-EB. In particular, GMS-DH as well as GMS-EB without augmentation are significantly faster than LKH and HGS.
For the purpose of ablation, apart from the previously mentioned benchmarks, we also evaluate a variant of GMS-DH with Pre-Pruning (GMS-DH PP) applied before the encoder in the MGMOTSP and MGMOCVRP settings. This pre-pruning is the same as for MBM, and the selection head is removed accordingly.
@@ 5.2 Zero-Shot Generalization
We also test our models on larger instances not seen during training, and report the results in Table 2. Both GMS-DH and GMS-EB remain competitive across both problems and all distributions in this zero-shot setting. Note that we refrain from testing the MOEAs here, as they already underperform significantly for smaller instances in Table 1. Moreover, MatNet is not applicable since it is limited by its embedding dimension (set to 128 in our case).
5.3 A Difficult Multigraph Problem
Finally, we revisit one of the main motivations behind this work: routing for problems without clear a priori edge selection. Thus, we evaluate the performance on a more complex problem. Specifically, inspired by Ben Ticha et al. (2017), we consider an MGMOTSP with Time-Windows (MGMOTSPTW). Each edge is associated with a time and distance, while the objectives correspond to the number of violated time-windows and the total distance. Remark that, unlike the MGMOTSP and MGMOCVRP, it is not clear in this case which edges can be pruned due to the more intricate time-window objective.
@@ 6 Conclusion
In this paper, we propose two learning-based approaches for solving multi-objective routing problems, such as the MOTSP and MOCVRP, on multigraphs and asymmetric graphs. Both methods utilize graph neural networks for encoding the graph structure but differ in their route construction strategies. The first method is edge-based and directly builds tours on the multigraph, while the second employs a two-stage decoding process, consisting of non-autoregressive pruning of the multigraph followed by route construction.
Empirical results show that both methods perform well across a variety of settings. While the edge-based model is simpler and usually demonstrates slightly superior performance, the second model is more scalable due to lower memory and runtime demands. In the future, we will explore alternative decomposition methods to pruning, to potentially obtain even more scalable methods.
@@ Acknowledgements
Appendix A Extended Related Work
A.1 Single-Objective Learning for Routing
In the non-autoregressive category, some works are Joshi et al. (2019); Fu et al. (2021); Qiu et al. (2022); Sun and Yang (2023). Typically, these methods require more sophisticated procedures than simple sampling in the decoding, such as Monte–Carlo tree search, to perform competitively, but scale better to large instances due to lower space and time complexity. Thus, an interesting recent work combining autoregressive and non-autoregressive construction to solve large instances (up to $10^{5}$ nodes for the TSP) while maintaining performance is Ye et al. (2024).
Besides constructive approaches, improvement-based methods usually obtain great results, but with longer running-times, by refining initial solutions. Some examples are Chen and Tian (2019); Lu et al. (2020); Hottung and Tierney (2020); Hudson et al. (2022).
A.2 Multi-Objective Learning for Routing
Early works treating MO routing using learning utilize multiple models, where one model is responsible for one preference between the objectives (Li et al., 2021; Wu et al., 2020). This quickly becomes infeasible as the number of models required grows exponentially with the number of objectives (Ehrgott, 2005). Pareto-set learning for multi-objective routing using a single model was proposed by Lin et al. (2022). They used the hyper-networks of Navon et al. (2021) to condition the decoder on the preference, which ensures the encoder only needs to be run once per instance. Recent studies that feature improved results or extend to more intricate VRP variants with a similar conditioning method include Li et al. (2023), Lu et al. (2024), Wu et al. (2024), Wu et al. (2025) and Fan et al. (2025a). An alternative approach is to use meta-learning to train a single model that can be quickly fine-tuned over a few steps to solve individual subproblems. Two representative works in this direction are Zhang et al. (2021) and Chen et al. (2023a).
Notably, all these prior methods utilize transformer-based architectures and operate on node coordinates, ensuring they are not suitable for asymmetric problems or multigraph problems. Besides tackling asymmetric problems, our work is novel in the MO routing setting as most methods utilize autoregressive construction, while one of our models combine autoregressive and non-autoregressive construction.
Another significant recent work is that of Chen et al. (2023b), in which two diversity enhancement methods, together labeled NHDE, are introduced. The authors combine NHDE with both the meta-learning approach and hyper-network approach to achieve improved results. Similarly, Fan et al. (2025b) develop a model-agnostic framework, POCCO, for improved training using pairwise preference learning and conditional computation blocks. Integrating GMS with NHDE or POCCO is possible in principle, but outside the scope of this paper.
A.3 Routing on Multigraphs
Appendix B Extended Motivation
In this section, we further motivate our work by explaining why most existing methods are unsuitable for routing on multigraphs. We also explain why transforming the underlying multigraph and applying vanilla node-based construction has significant scaling issues and other disadvantages. Finally, we discuss the technical novelty of the various components in GMS-DH and GMS-EB.
B.1 Neural Routing Methods Extended to Multigraphs
There exists many single-objective neural methods for routing capable of encoding simple asymmetric graphs (Kwon et al., 2021; Sun and Yang, 2023; Lischka et al., 2025; Drakulic et al., 2025). However, a significant amount of these, in particular transformer-based architectures like MatNet (Kwon et al., 2021) and GOAL (Drakulic et al., 2025), are unsuitable for encoding multigraphs. Both use distance-matrices as input to mixed attention blocks. Representing multigraphs through distance matrices ($N\times N\times D$ dimensions) by e.g., concatenation along the feature dimension will result in violated permutation invariance among parallel edges, which is key to maintain. As such, to extend these transformer-based encoders to multigraph inputs, one would need to design edge-permutation invariant attention layers, which to our knowledge has not been attempted. Note that we apply MatNet to multigraphs by pre-pruning so that there is only one edge between each node pair. This works, but learned pruning after encoding is more suitable in the general case when good pruning heuristics might not be available.
Additionally, some recent methods (Lischka et al., 2025; Meng et al., 2025) are edge-based and can encode multigraphs, but rely on node-selection in the decoding, ensuring they are not applicable to routing on multigraphs. An even smaller subset of related work, e.g., Zhou et al. (2025) in the autoregressive paradigm and Sun and Yang (2023) in the non-autoregressive paradigm, provide edge selections in the decoding, which would in theory make them applicable to multigraph routing. However this extension has not been attempted in any previous work we are aware of, and we hypothesize that direct applications would lead to problems akin to GMS-EB (which can be viewed as an initial extension of GREAT (Lischka et al., 2025)), where the scaling is slow due to not explicitly designing for the multigraph structure.
B.2 Node-Based Construction and Graph Transformations